# Airdrops: Giving Money Away Is Harder Than It Seems

### v1

@@ 3.2. Token Distribution Graphs
(a) Absolute number of unique daily active addresses.
(b) Daily active addresses, relative to the average for the 50 days before Arbitrum’s airdrop.
@@
Figure 3. Unique daily active addresses per platform. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
@@
Figure 4. Median daily transaction fees, in USD.
4. Measuring the Airdrop Lift
@@
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic. In this section, we examine this through the lens of some relevant metrics applied to the Arbitrum airdrop. We rely on data source from Growthepie (Growthepie, 2023) and list details in Section A in appendix.
4.1. Unique Active Addresses
@@
Figure 3 examines the number of unique addresses that were active on each platform using two different approaches. The Figure 2(a) depicts the absolute number of addresses. The Figure 2(b) shows the relative amount of unique daily active addresses for each platform, when compared to the same platform’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 (Rustgi, [n. d.]). This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users (Barthere, 2023).
@@
Observation 1: others outperformed Arbitrum, without an airdrop. The absolute number of addresses has increased for Arbitrum post-airdrop, and remained at least 50% higher than its pre-airdrop level in the four months since then. But, by looking at other platforms, it becomes apparent that a similar result can be attained without an airdrop. This is demonstrated by Optimism’s trajectory in Figure 2(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 (Nijkerk, 2023). This is further exemplified by zkSync Era overtaking Arbitrum’s in Figure 2(a), two months after the airdrop.
@@
Observation 2: fees may explain the closing bap between Arbitrum and Optimism. Arbitrum leads over Optimism with respect to the number of unique daily active addresses. While in the 50 days before the airdrop Arbitrum has enjoyed an average of 2.6 times more unique daily active addresses than Optimism, this gap has recently started to close: the average over the last 50 days in our dataset is considerably lower, and is equal to 1.83 times more. Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 4.
@@
Observation 3: the number of unique addresses oscillates over time. Some of the previous observations exhibit oscillatory behavior, in particular in Figure 2(b) depicting the relative amount of unique addresses: rising, reaching a peak, and then subsequently dropping, all in rapid succession. In particular, the phases of Optimism and Arbitrum are inverted regarding the relative amount of unique addresses. One may attempt to explain this by reasoning that when a peak is reached in one of the platforms, fees should increase due to the load, thus prompting cost-conscious users to turn to a different platform. But, when considering the median transaction fees paid per day (per Figure 4), it is apparent that fees do not have corresponding peaks and dips. In fact, Arbitrum consistently had lower median fees mid-May.
Caveat: the unique address metric is gameable. The overall picture painted so far is that with regard to the daily unique account metric, Arbitrum’s airdrop did not succeed in generating long-lasting user involvement. An important caveat of this metric is that in blockchains, users can typically costlessly create and use multiple addresses. This practice is also called a sybil attack, with attackers also known as sybils. Sybil attacks may benefit users, as some airdrops limit the amount of tokens that a single address can receive. Thus, using the number of unique addresses as a metric may paint an incomplete picture: the large changes we observed may be caused by sybil attackers. Although some employ sophisticated tools to identify sybil addresses, these are heuristic in nature and may fail to capture all of them. Thus, we need to consider other metrics that are hopefully more sybil-resistant and representative of genuine user activity.
4.2. Transaction-related Metrics
@@
A natural next step is to examine transaction-related metrics. In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee (Gafni and Yaish, 2022), when ignoring retroactive rebates or the possibility of platforms operating at a loss (Optimism, 2023).
Observation 4: the Arbitrum-Optimism gap is smaller, when considering transactions. Figure 4(a) depicts the absolute daily transaction count per platform, and Figure 4(b) shows the relative amount of transaction count for each platform compared to the same platform’s average transaction count for the period between February 1st and March 22nd, 2023. These figures are quite similar to the corresponding ones for the unique address count metric.
@@
Some notable differences are that with respect to transaction counts, the gap between Arbitrum and Optimism has nearly closed entirely by the end of July. Furthermore, Figure 4(b) shows that the daily transaction count for Immutable X has almost halved since Arbitrum’s airdrop. Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the platform.
@@
(b) Daily transactions relative to the average for the 50 days before Arbitrum’s airdrop.
Figure 5. Daily number of transactions per platform. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
Observation 5: Arbitrum’s per address transaction count dropped since the airdrop. To better understand how Arbitrum’s airdrop affected user engagement in other platforms, we now present the Figure 6, which depicts the relative average daily transaction count per unique address. The figure shows that since the drop, Arbitrum’s per-user transaction count has fallen to less than 75% of its pre-airdrop level.
Figure 6. Daily transactions per address, relative to the average before Arbitrum’s airdrop.
Caveat: transaction counts are inaccurate when ignoring fees. Relying on the previous metrics may be inaccurate: a high transaction count does not necessarily signal high user engagement. In particular, some platforms require users to perform a certain number of transactions over some time frame to become eligible for their airdrops. Thus, if fees are low, a high number of transactions per account may be caused by users who perform economically-useless transactions, for the sole purpose of receiving the drop. This brings to mind Goodhart’s law (Manheim and Garrabrant, 2019), which posits that “when a measure becomes a target, it ceases to be a good measure”.
Observation 6: since June, average transaction fees are similar for all platforms. Conceptually, good metrics should measure a user’s amount of skin in the game. A possible proxy for that is the amount of funds the user is willing to expend for interacting with the platform. When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all platforms are quite similar in both regards. Furthermore, when comparing the average fee per transaction to the median one (per Figures 4 and 7), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
Figure 7. Daily average transaction fee spent per transaction per day, in USD.
Another interesting metric is the relative average fees per address, compared to the average in the 50 days preceding Arbitrum’s airdrop, shown in Figure 8. One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other platforms.
Figure 8. Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
Figure 9. Total fees relative to the average for the 50 days before Arbitrum’s airdrop.
Observation 7: Arbitrum’s total daily fees spiked around the time of the airdrop. Figure 9 shows that Arbitrum’s airdrop did not confer a substantial advantage over its competitors when it comes to its income stream from transaction fees. Although the total amount of fees it earned on the day of the drop reached a peak, this was momentary.
Indeed, at the relative lead, Arbitrum held over Optimism with respect to this metric, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
Figure 10. Daily TVL increase relative to the average for the 50 days before Arbitrum’s airdrop.
4.3. Total Value Locked
The Total Value Locked (TVL) of a platform is a metric that measure the combined value of all assets stored in that platform.
@@ 4.3. Total Value Locked
5. Common Pitfalls
@@
Airdrops share similarities with traditional loyalty programs, such as newcomer bonuses offered to new customers by banks and credit card companies. Furthermore, they also share some common pitfalls. The blockchain setting and the mechanics adopted by most airdrops may exacerbate these pitfalls, or introduce new ones.
@@
Airdrop farmers. Sophisticated users known as airdrop farmers employ elaborate tactics to maximize the amounts of airdrop tokens they receive. Blockchain platforms have adopted various measures to curb the possibility of airdrop farmers gaming the system and obtaining an inappropriate amount of rewards. For example, a popular approach is to limit the amount of rewards that a single user can receive. Alas, a single user can pose as multiple users, in a move called a sybil attack.
@@
Thus, platforms resort to Proof-of-Humanity services, such as Gitcoin passport. These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human. Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH (Gitcoin, 2023). Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop (Barthere, 2023), but this may lead to false negatives (X-explore, 2023). There are some recent examples of platforms like Celestia considering using sybil resistance solutions (Labs, 2023c).
There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions (Protocol, 2023), to sharing posts on social media (Labs, 2023b). These tasks sometimes appear arbitrary, leading to frustration among users.
@@
The problem with the aforementioned methods is that they are easy to automate (Jumper-sketch, 2023) and cheaply fooled, even more-so when platforms offer retroactive rebates on transaction fees (Optimism, 2023). Furthermore, the reliance of many platforms on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops. Even using biometric identification does not guarantee sybil-proofness (TrustaLabs, 2023). A different approach used by platforms is to announce airdrops that retroactively reward users that were active before the announcement. But, farmers can prepare in advance by interacting with these platforms even when no airdrop is officially announced (Copeland, 2023), as was the case in dYdX’s airdrop (Edwards, 2022).
@@
Distributing governance tokens may be dangerous. Some platforms airdrop governance tokens, as a means to decentralize their governance process. Such tokens allow holders to vote in the platforms’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value (Dotan et al., 2023).
@@
There is some empirical evidence that suggests that airdrops of governance tokens may be more successful than airdrops of non-governance tokens. In a recent analysis of the market capitalization of crypto tokens performed by Makridis et al. (Makridis et al., 2023), the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped. But, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
@@
Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system (Messias et al., 2023a; Feichtinger et al., 2023; Sharma et al., 2023). Also, to users that do not have the platform’s best interests in mind, and may even act to the determent of the platform.
6. Design Guidelines
@@
The pitfalls outlined earlier, while indeed sobering, may inform platforms embarking on the airdrop journey, and can shed light on the path to potential success.
@@
Alternative incentives may better sustain user engagement. An airdrop’s long-term benefit to a platform can be tricky to measure, given that the potential benefits may be indirect and hard to quantify in advance, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from platforms to users and is thereafter no longer available for other uses. Furthermore, some costs and effects cannot be easily foreseen, like those arising from distributing governance tokens to users.
@@
Instead of using airdrops, a community could adopt other measures which have a more predictable relation between costs and benefits. For example, a very natural alternative to airdrops is for the community to vote to programmatically reward power users with discounts on future interactions. When considering L2s, these discounts can be given on transaction fees. Thus, if a user receives this reward, it necessarily has to use the platform again in the future to benefit from the incentive, thus effectively fostering continued user engagement. Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the platform, and the platform’s costs go solely towards users actively using the system.
@@
Nevertheless, while fee discounts are more focused towards incentivizing continued interaction with the platform, they are not without their own challenges. For example, one should carefully design a proper discount mechanism, covering both the eligibility criteria for users, and the height of awarded discounts. Furthermore, it is not immediately apparent that discounts can attract users to a similar degree as the more immediate and tangible rewards offered by airdrops. Another option is to perform multiple airdrops over a relatively long period of time, instead of a single one.

### v2

@@ 3.2 Token Distribution Graphs
(a) Absolute number.
(b) Relative daily active addresses.
@@
Figure 4: Unique daily active addresses per protocol: (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
Figure 5: Ratio between the number of unique daily active addresses in Arbitrum and Optimism.
(a) Daily median fees.
(b) Daily average fees.
(c) Average fees spent per unique address per day.
@@
Figure 6: Daily transaction fees in USD. (a) Median; (b) Average; and (c) Per unique address.
4 Measuring the Airdrop Lift
@@
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [21]. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [30, 21]. In this section, we examine this through the lens of some relevant metrics applied to the Arbitrum airdrop. We rely on data source from Growthepie [34]. This platform supplies various metrics, including daily transaction count, daily active addresses, median transaction cost, TVL, fees paid by users, and stablecoin market cap, for several key protocol such as zkSync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM. To enhance clarity, all these metrics were subjected to a 7-day moving average smoothing process.
4.1 Unique Active Addresses
@@
Figure 4 examines the number of unique addresses that were active on each protocol using two different approaches. The Figure 4(a) depicts the absolute number of addresses. The Figure 4(b) shows the relative amount of unique daily active addresses for each protocol, when compared to the same protocol’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 [66]. This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users [7].
{observation}
@@
others outperformed Arbitrum, without an airdrop. The absolute number of addresses has increased for Arbitrum post-airdrop, and remained at least 50% higher than its pre-airdrop level in the four months since then. However, by looking at other protocols, it becomes apparent that a similar result can be attained without an airdrop. This is demonstrated by Optimism’s trajectory in Figure 4(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 [62]. This is further exemplified by zkSync Era overtaking Arbitrum’s in Figure 4(a), two months after the airdrop.
{observation}
@@
fees may explain the closing gap between Arbitrum and Optimism. Arbitrum leads over Optimism with respect to the number of unique daily active addresses. We turn our focus to Figure 5, which depicts Arbitrum’s lead over Optimism with respect to the number of unique daily active addresses. While in the 50 days before the airdrop Arbitrum has enjoyed an average of 2.6 times more unique daily active addresses than Optimism, this gap has recently started to close: the average over the last 50 days in our dataset is considerably lower, and is equal to 1.83 times more. Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 6(a). We note that with respect to the average fee, the gap between the protocols is even smaller over the same period as shown in Figure 6(b).
{observation}
@@
the number of unique addresses oscillates over time. Some of the previous observations exhibit oscillatory behavior, in particular in Figure 4(b) depicting the relative amount of unique addresses: rising, reaching a peak, and then subsequently dropping, all in rapid succession. In particular, the phases of Optimism and Arbitrum are inverted regarding the relative amount of unique addresses. One may attempt to explain this by reasoning that when a peak is reached in one of the protocols, fees should increase due to the load, thus prompting cost-conscious users to turn to a different protocols. But, when considering the median transaction fees paid per day (per Figure 6(a)), it is apparent that fees do not have corresponding peaks and dips. In fact, Arbitrum consistently had lower median fees in mid-May.
[…]
4.2 Transaction-related Metrics
@@
A natural next step is to examine transaction-related metrics. In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee [32], when ignoring retroactive rebates or the possibility of blockchain protocols operating at a loss [63].
{observation}
[…]
@@
Some notable differences are that with respect to transaction counts, the gap between Arbitrum and Optimism has nearly closed entirely by the end of July. Furthermore, Figure 7(b) shows that the daily transaction count for Immutable X has almost halved since Arbitrum’s airdrop. Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the protocol.
@@
(b) Relative daily transactions.
Figure 7: Daily number of transactions per protocol. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
{observation}
Arbitrum’s per address transaction count dropped since the airdrop. To better understand how Arbitrum’s airdrop affected user engagement in other protocols, we now present the Figure 8. Figure 8(b), which depicts the relative average daily transaction count per unique address. The figure shows that since the drop, Arbitrum’s per-user transaction count has fallen to less than 75% of its pre-airdrop level.
(a) Average number.
(b) Relative to average before Arbitrum’s airdrop.
Figure 8: Daily transactions per address.
@@
{observation}
since June, average transaction fees are similar for all protocols. Conceptually, good metrics should measure a user’s amount of skin in the game. A possible proxy for that is the amount of funds the user is willing to expend for interacting with the protocol. When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all protocols are quite similar in both regards. Furthermore, when comparing the average fee per transaction to the median one (per Figures 6(a) and 6(b)), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
Another interesting metric is the relative average fees per address, compared to the average in the 50 days preceding Arbitrum’s airdrop, shown in Figure 9. One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other protocols.
Figure 9: Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
(a) Total daily transaction fees (USD).
(b) Relative to average before Arbitrum’s airdrop.
Figure 10: Total fees before and after the Arbitrum airdrop.
Figure 11: Ratio between the daily fees spent in Arbitrum and Optimism.
{observation}
Arbitrum’s total daily fees spiked around the time of the airdrop. Figure 10 shows that Arbitrum’s airdrop did not confer a substantial advantage over its competitors when it comes to its income stream from transaction fees. Although the total amount of fees it earned on the day of the drop reached a peak, this was momentary.
Indeed, per Figure 11, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
(a) Daily TVL, in USD.
(b) Relative to the average prior Arbitrum’s airdrop.
Figure 12: Total Value Locked (TVL) growth: (a) Daily TVL in USD; and (b) relative to the average for the 50 days before Arbitrum’s airdrop.
4.3 Total Value Locked
The Total Value Locked (TVL) of a protocol is a metric that measure the combined value of all assets stored in that protocol.
{observation}
[…]
5 Common Airdrop Design Challenges
@@
Airdrops share similarities with traditional loyalty programs, such as newcomer bonuses offered to new customers by banks and credit card companies. Furthermore, they also share some common design challenges. The blockchain setting and the mechanics adopted by most airdrops may exacerbate these pitfalls, or introduce new ones.
5.1 Airdrop Farmers
@@
Sophisticated users known as airdrop farmers employ elaborate tactics to maximize the amounts of airdrop tokens they receive. Blockchain protocols have adopted various disproportionate measures to curb the possibility of airdrop farmers gaming the system. For example, a popular approach is to limit the amount of rewards that a single user can receive.
@@
Thus, protocols resort to Proof-of-Humanity services, such as Gitcoin passport [33]. These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human. Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH [33]. Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop [7], but this may lead to false negatives [76]. There are some recent examples of protocols like Celestia considering using sybil resistance solutions [47].
@@
There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions [65], to sharing posts on social media [45]. These tasks sometimes appear arbitrary, leading to frustration among users. The problem with the aforementioned methods is that they are easy to automate [40] and cheaply fooled, even more-so when protocols offer retroactive rebates on transaction fees [63]. Furthermore, the reliance of many protocols on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops, even with biometric identification, full sybil resistance is not guaranteed [72]. A different approach used by protocols is to announce airdrops that retroactively reward users that were active before the announcement. Nevertheless, farmers can prepare in advance by interacting with these protocols even when no airdrop has officially been announced [14], as was the case in dYdX’s airdrop [18].
@@ 5.2 Decentralized Governance Threats
Some protocols airdrop governance tokens, as a means to decentralize their governance process. However, distributing governance tokens may be dangerous. Such tokens allow holders to vote in the protocols’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value [17].
@@
There is some empirical evidence that suggests that airdrops of governance tokens may be more successful than airdrops of non-governance tokens. In a recent analysis of the market capitalization of crypto tokens performed by Makridis et al. [53], the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped. However, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
@@
Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system [56, 22, 68]. Also, to users that do not have the protocol’s best interests in mind, and may even act to the determent of the protocol.
5.3 Insider Trading
It is a problematic occurrence within organizations where individuals exploit privileged information to gain a financial advantage over other users of the protocol. This practice is widely recognized as a violation of securities laws in traditional financial markets [6], and engaging in such activities can lead to negative responses from the blockchain community [69, 61, 23].
When someone within a protocol utilizes privileged information to enhance their profits during an airdrop, it can trigger a backlash from the blockchain community. Insider traders have advanced knowledge of the metrics used to determine eligibility and the amount of an airdrop for each address. This knowledge empowers them to manipulate the system and make a profit. For example, there were claims that a head of growth at AltLayer might have used insider information to make a $200000$ USD profit from their airdrop, although later dismissed as a mere coincidence [69]. However, such incidents can undermine the trust users place in these protocols.
This issue also raises concerns about fairness, as some users have access to superior and more accurate information than others. Identifying these traders is a challenging problem, making it crucial for protocols to provide extensive information to their users. In addition, incentives for blockchain data analytics companies and research groups to conduct post-airdrop data audits can help identify insider traders by analyzing address details and transfer patterns [23]. For this to happen, data availability is key. Therefore, protocols should ensure transparency and encourage thorough analysis to maintain integrity and fairness within the blockchain community.
6 Design Guidelines
@@
The airdrop design challenges outlined earlier, while indeed sobering, may inform blockchain protocols embarking on the airdrop journey, and can shed light on the path to potential success.
6.1 Alternative Incentives may Better Sustain User Engagement
@@
An airdrop’s long-term benefit to a protocol can be tricky to measure, given that the potential benefits may be indirect and hard to quantify, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from protocols to users and is thereafter no longer available for other uses. Furthermore, some costs and effects cannot be easily foreseen, like those arising from distributing governance tokens to users.
@@
Instead of using airdrops, a community could adopt other measures which have a more predictable relation between costs and benefits. For example, a very natural alternative to airdrops is for the community to vote to programmatically reward power users with discounts on future interactions. When considering L2s, these discounts can be given on transaction fees. Thus, if a user receives this reward, it necessarily has to use the protocol again in the future to benefit from the incentive, thus effectively fostering continued user engagement. Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the protocol, and the protocol’s costs go solely towards users actively using the system.
@@
Nevertheless, while fee discounts are more focused towards incentivizing continued interaction within the protocol, they are not without their own challenges. For example, one should carefully design a proper discount mechanism, covering both the eligibility criteria for users, and the height of awarded discounts. Furthermore, it is not immediately apparent that discounts can attract users to a similar degree as the more immediate and tangible rewards offered by airdrops. Another option is to perform multiple airdrops over a relatively long period of time, instead of a single one. Although this is still susceptible to some of the pitfalls of standard one-shot airdrops, it may serve to ensure the long-term engagement of the community, and thus avoid the decline in user adoption that some experience immediately after an airdrop.
A more innovative approach was adopted by Blast, which introduced a point-based reward program [55]. Under this program, users receive rewards based on the points they accumulate through various activities, such as bridging tokens to the protocol (i.e., transferring users’ funds from another protocol) and participating in a referral program where bringing in more users increases their points. Interestingly, Blast received 1.1 billion USD in deposits even before its official launch [43]. This method provides an evaluative metric for users’ contributions to the protocol, structured around a referral-program model.
Furthermore, the incorporation of innovative distribution mechanisms demonstrates as a critical role in mitigating the presence of sybils within whitelisted addresses eligible for an airdrop. For example, Celestia has proposed a distinctive design that employs GitHub commits as a proxy to assess a user’s contribution to the blockchain ecosystem [27]. Nevertheless, a potential concern emerges in regard to the possibility of users or farmers engaging in the generation of counterfeit activities on GitHub, aiming to exploit other protocols’ airdrops adopting a strategy analogous to Celestia’s. Among other metrics, would be to focus on metrics that are resistant to programmable manipulation, adding a layer of difficulty or cost to automated user accounts.
6.2 Target Well-known Reputable Entities